OpenStudy (sleepyjess):

A projectile is thrown upward so that its distance above the ground after t seconds is \(h = - 16t^2 + 480t\). After how many seconds does it reach its maximum height? 15 s 22.5 s 7 s 30 s

OpenStudy (sleepyjess):

@hartnn last one i hope

OpenStudy (sleepyjess):

I have done these before but I can't remember how I did it and I can't find my notes on it either

hartnn (hartnn):

for \(y= ax^2+bx+c\) the maximum occurs at \(-b/2a\)

hartnn (hartnn):

h=−16t^2+480t. compare this with h =at^2 +bt+c find a,b then use that formula

OpenStudy (sleepyjess):

so -480/-32 15

hartnn (hartnn):

yes :)

OpenStudy (sleepyjess):

Thank you so much for all of your help!

hartnn (hartnn):

welcome ^_^

OpenStudy (nincompoop):

there's another way to solve this, but that's probably for another day or semester

OpenStudy (nincompoop):

it is to make sense why -b/2a is legitimate

OpenStudy (javk):

After 15s You could also resolve this by imagining its graph Initially as time increases height also increases Once maximum height is reached, with any increase in time height decreases Therefore the graph has zero gradient at maximum height By differentiation we get: \[dh/dt=32t+480\] Where gradient equals zero, therefore dh/dt=0 \[32t+480=0\] \[t=15\]